OpenStudy (anonymous):

HELP PLEASE solve the equation: -2/x+4 = 4/x+3 asap!

OpenStudy (anonymous):

i will medal anyone who can help

OpenStudy (freckles):

\[\frac{-2}{x}+4=\frac{4}{x}+3 ?\]

OpenStudy (freckles):

I think you mean though -2/(x+4)=4/(x+3) which is: \[\frac{-2}{x+4}=\frac{4}{x+3}\]

OpenStudy (freckles):

if so multiply both sides by (x+3) and (x+4) you should wind up with -2(x+3)=4(x+4)

OpenStudy (freckles):

try solving that and let me know what you get

OpenStudy (anonymous):

-11/3?? @freckes

OpenStudy (anonymous):

and ya that is what i meant

OpenStudy (freckles):

-2x-6=4x+16 add 2x on both sides -6=6x+16 subtract 16 on both sides -22=6x divide both sides by 6 -22/6=x reduce -22=2(-11) and 6=2(3) so yes x=-11/3